RC meeting and Run Plan, Thursday 11th April

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 3:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  1. Completed yesterday:
    • Q2 = 5.5, W=3.02, x=0.40, low epsilon, SHMS =8.5 degrees, 70uA. 72 hours of running. ~850 Lambdas were collected (less than 10 per hour) by 1pm.
    • Peter's SIDIS high Q2 study.
    • SHMS angle change to 6.2 degrees completed at 5pm.
  2. SHMS dipole caused 10 hours lost beam time yesterday owl shift.
  3. Currently running the Q2 = 5.5, W = 3.02, x = 0.40, SHMS = 6.2 degrees: 10% completed
  4. Beam:
    • Unstable over night, PSS drop caused 4 hours down time this morning (9am-1pm).

Current Run Plan

  • We stay at this setting until Monday beam off
  • Data at 50uA and dummy target (2.5 hours) at 40uA.

Notices in Effect for Shift Crew

  • Please take a screenshot of the spectrometer angles when they are moved to a new position and make a log entry. If working on a kinematic that lasts several days take an image daily.
  • Pay attention to the notices posted in the counting house regarding HMS Q1/Q2.
  • If HMS interlocks when ramping, check the live camera image of the HMS power supplies for smoke/water before clearing the interlock.
  • Hit enter once you type in the angle when rotating(!)

Issue

  • Wifi connection in the hall does not work for Steve and Mike's laptop, but works for Mike's ipad. Rounter reset?

Special Requests